# Polynomial time sampling from log-smooth distributions in fixed dimension under semi-log-concavity of the forward diffusion with application to strongly dissipative distributions

### v2

@@ 3.3 Assembling the pieces
where $\lesssim$ hides a universal constant as well as log factors with respect to $K_{1},K_{2}$.
The proof is left in Appendix A.9 and gives the appropriate set of parameters to use provided $L,R,m_{2}(\mu),\beta$ are known. We also provide in Appendix a version of this theorem that handles the case where the constants $L,R,\beta,m_{2}(\mu)$ are unknown.
Remark 12
As the proof shows, one could have relaxed the uniform-in-time upper-bound on $\nabla^{2}\log(p_{t}^{V})$, $\nabla^{2}\log(p_{t}^{2V})$ into a time-dependent version. This relaxation would simply have replaced the term $\beta$ in our bounds with $\max_{1\leq k\leq N}\beta_{t_{k}}$.
4 Application to smooth and strongly dissipative potentials
In this section, we prove that the results above apply in the strong dissipative case.
Assumption 3 (Strong Dissipativity)
The potential $V$ is $\alpha$-strongly convex outside some ball of radius $R$ centered at $m\in\mathbb{R}^{d}$ denoted $B_{R}(m)$.
We prove that if Assumptions 2-3 hold, then Assumption 1 holds with $\beta\asymp L^{2}(R^{2}(L/\alpha)^{2}+d/\alpha)$. The proof is decomposed into two time regimes: first, when $t$ is sufficiently small, the following result ensures that the forward process remains log-smooth, and a fortiori semi-log-concave, when $\mu$ is.
Proposition 13 (Chen et al., 2023a [10])
If $V$ is $C^{2}$ and $L$-smooth and if $t$ is such that $1-e^{-2t}\leq\frac{e^{-t}}{L}$, then it holds that $\|\nabla^{2}\log(p_{t}^{V})\|\leq 2Le^{t}$.
Hence, we only need to control $\nabla^{2}\log(p_{t}^{V})$ when $t\gtrsim 1$. Tweedie’s formula (see Appendix A.1) yields the identity $\nabla^{2}\log(p_{t}^{V})(x)-\nabla^{2}\log(\pi)(x)=\frac{e^{-2t}}{1-e^{-2t}}\left(\frac{\text{Cov}_{q_{t,x}}(X)}{1-e^{-2t}}-I_{d}\right)\,,$ where $q_{t,x}\propto e^{-V(z)}e^{-\frac{e^{-2t}\|z-xe^{t}\|^{2}}{2(1-e^{-2t})}}$. Hence, in order to control $\nabla^{2}\log(p_{t}^{V})$ from above, it suffices to upper-bound $\text{Cov}_{q_{t,x}}$ independently of $x$. Now remark that if $V$ is $\alpha$-strongly convex outside some ball of radius $R$ with condition number $\kappa\coloneqq L/\alpha$, then $q_{t,x}$ remains at least $\alpha$-strongly log-concave outside the same ball of radius $R$ and has most condition number $\kappa$. We thus derived the following result which upper-bounds the variance (and a fortiori the covariance matrix) of log-smooth and strongly log-concave measures outside some ball of radius $R$.

### v3

@@ 5 Sketch of proof of the main result
where $\lesssim$ hides a universal constant as well as log factors with respect to $d,\epsilon^{-1},a,b,\beta$. In particular, the error above is achieved in $\sum_{k=1}^{N}n_{k}=d\epsilon^{-2(d+2)}$ queries to $V$.
The proof is deferred to Appendix F and is mainly an application of the results collected above.
6 Conclusion
In this article, we successfully applied the reduction from sampling to intermediate score estimation, initiated over the past three years by Chen et al. [2023b, a], Conforti et al. [2025], Benton et al. [2024], to the problem of low dimensional multi-modal sampling. Using the self-normalized estimator of the scores, our results provide polynomial query complexity guarantees in fixed dimension, apply to general Gaussian mixtures, and do not require prior knowledge of the target distribution’s constants. Interesting future directions include extending theoretical guarantees to more general multi-modal distributions with heavy tails for instance.
We note that our sampling algorithm is based on time-reversed diffusions, which have recently gained traction for sampling from unnormalized densities. Our method stands out by offering rigorous, non-asymptotic theoretical guarantees on query complexity, especially in the presence of score estimation error that we precisely quantify. Such theoretical guarantees are scarce and we believe our results are therefore a meaningful and timely contribution to the field.
7 Acknowledgements
This work was supported by the Agence Nationale de la Recherche (ANR) through the JCJC WOS project and the PEPR PDE-AI project (ANR-23-PEIA-0004). We thank Pierre Monmarché for his help in proving Lemma 13.
Appendix A Expressions the intermediate scores
We next provide derivations of the Monte Carlo estimators of the intermediate scores discussed in section 4.2.
The probability law along the reverse diffusion path has density
$p_{t}^{V}(x)$ $=\frac{1}{\sqrt{1-\lambda_{t}}}\pi\left(\frac{x}{\sqrt{1-\lambda_{t}}}\right)\ast\frac{1}{\sqrt{\lambda_{t}}}\mu\left(\frac{x}{\sqrt{\lambda_{t}}}\right)$ (6)
where $\pi$ is a standard Gaussian and $\mu(\cdot)\propto\exp(-V(\cdot))$ is the target distribution
Case 1: write the convolution as an integral against the proposal distribution
We have
$p_{t}^{V}(x)$ $=\frac{1}{Z_{\lambda_{t}}}\int_{y\in\mathbb{R}^{d}}\mu\left(\frac{x-y}{\sqrt{\lambda_{t}}}\right)\pi\left(\frac{y}{\sqrt{1-\lambda_{t}}}\right)dy=\frac{1}{Z_{\lambda_{t}}}\int f(x,y)dy$ (7)
where we denoted the integrand by $f(x,y)$. We can compute the score
$\nabla\log p_{t}^{V}(x)$ $=\frac{\nabla p_{t}^{V}(x)}{p_{t}^{V}(x)}=\frac{\int\nabla f(x,y)dy}{\int f(x,y)dy}=\frac{1}{\sqrt{\lambda_{t}}}\int\nabla\log\mu\left(\frac{x-y}{\sqrt{\lambda_{t}}}\right)\frac{f(x,y)}{\int f(x,y)dy}dy\kern 5.0pt.$ (8)
From this, we can either define a Monte Carlo estimator as
$\nabla\log p_{t}^{V}(x)$ $=\frac{1}{\sqrt{\lambda_{t}}}\mathbb{E}_{y|x}\left[\nabla\log\mu\left(\frac{x-y}{\sqrt{\lambda_{t}}}\right)\right],\quad y|x\sim\propto f(x,y)$ (9)
where $f(x,y)$ is a smoothened version of the target distribution distribution. Replacing $\pi$ with a standard Gaussian. Or else, by unpacking $f(x,y)$ and using the proposal as the sampling distribution as the integration variable $y$ appears in it,
$\nabla\log p_{t}^{V}(x)$ $=\frac{1}{\sqrt{\lambda_{t}}}\frac{\mathbb{E}_{y}\left[\nabla\log\mu\left(\frac{x-y}{\sqrt{\lambda_{t}}}\right)\mu\left(\frac{x-y}{\sqrt{\lambda_{t}}}\right)\right]}{\mathbb{E}_{y}[\mu\left(\frac{x-y}{\sqrt{\lambda_{t}}}\right)]},\quad y\sim\pi\left(\frac{y}{\sqrt{1-\lambda_{t}}}\right)$ (10)
Case 2: write the convolution as an integral against the target distribution
We now have
$p_{t}^{V}(x)$ $=\frac{1}{Z_{\lambda_{t}}}\int_{y\in\mathbb{R}^{d}}\mu\left(\frac{y}{\sqrt{\lambda_{t}}}\right)\pi\left(\frac{x-y}{\sqrt{1-\lambda_{t}}}\right)dy=\frac{1}{Z_{\lambda_{t}}}\int f(x,y)dy$ (11)
where we now denote the integrand by $f(x,y)$. We can compute the score
$\nabla\log p_{t}^{V}(x)$ $=\frac{\nabla p_{t}^{V}(x)}{p_{t}^{V}(x)}=\frac{\int\nabla f(x,y)dy}{\int f(x,y)dy}$ (12)
$=\frac{1}{\sqrt{1-\lambda_{t}}}\int\nabla\log\pi\left(\frac{x-y}{\sqrt{1-\lambda_{t}}}\right)\frac{f(x,y)}{\int f(x,y)dy}dy\kern 5.0pt.$ (13)
From this, we can either define a Monte Carlo estimator as
$\nabla\log p_{t}^{V}(x)$ $=\frac{1}{\sqrt{1-\lambda_{t}}}\mathbb{E}_{y|x}\left[\nabla\log\pi\left(\frac{x-y}{\sqrt{1-\lambda_{t}}}\right)\right],\quad y|x\sim\propto f(x,y)$ (14)
where $f(x,y)$ is a smoothened version of the target distribution distribution. Or else, by upacking $f(x,y)$ and using the target as the sampling distribution as the integration variable $y$ appears in it,
$\nabla\log p_{t}^{V}(x)$ $=\frac{1}{\sqrt{1-\lambda_{t}}}\frac{\mathbb{E}_{y}\left[\nabla\log\pi\left(\frac{x-y}{\sqrt{1-\lambda_{t}}}\right)\pi\left(\frac{x-y}{\sqrt{1-\lambda_{t}}}\right)\right]}{\mathbb{E}_{y}[\pi\left(\frac{x-y}{\sqrt{1-\lambda_{t}}}\right)]},\quad y\sim\mu\left(\frac{y}{\sqrt{\lambda_{t}}}\right)$ (15)
which is not very useful given that sampling the target is hard in the first place. However, we can use the proposal distribution as the sampling distribution as the integration variable $y$ appears in it as well. To see this, recall that
$\pi\left(\frac{x-y}{\sqrt{1-\lambda_{t}}}\right)$ $=\mathcal{N}\left(\frac{x-y}{\sqrt{1-\lambda_{t}}};0,I_{d}\right)=\sqrt{1-\lambda_{t}}\mathcal{N}\left(x-y;0,(1-\lambda_{t})I_{d}\right)$ (16)
$=\sqrt{1-\lambda_{t}}\mathcal{N}\left(y;x,(1-\lambda_{t})I_{d}\right)$ (17)
This leads to
$\nabla\log p_{t}^{V}(x)$ $=\frac{1}{1-\lambda_{t}}\frac{\mathbb{E}_{y}\left[(y-x)\mu\left(\frac{y}{\sqrt{\lambda_{t}}}\right)\right]}{\mathbb{E}_{y}\left[\mu\left(\frac{y}{\sqrt{\lambda_{t}}}\right)\right]},\quad y\sim\mathcal{N}(y;x,(1-\lambda_{t})I_{d})$ (18)
$=\frac{1}{\sqrt{1-\lambda_{t}}}\frac{\mathbb{E}_{z}\left[z\mu\left(\frac{x+\sqrt{1-\lambda_{t}}z}{\sqrt{\lambda_{t}}}\right)\right]}{\mathbb{E}_{z}\left[\mu\left(\frac{x+\sqrt{1-\lambda_{t}}z}{\sqrt{\lambda_{t}}}\right)\right]},\quad z\sim\mathcal{N}(z;0,I_{d}).$ (19)
Using a change of variables, we obtain
$\nabla\log(p_{t}^{V})(x)=\frac{-1}{1-\lambda_{t}}\frac{\mathbb{E}\big[Y_{t}\exp(-V(\frac{x-Y_{t}}{\sqrt{\lambda_{t}}}))\big]}{\mathbb{E}\big[\exp(-V(\frac{x-Y_{t}}{\sqrt{\lambda_{t}}}))\big]}$ (20)
with $Y_{t}\sim\mathcal{N}(0,(1-\lambda_{t})I_{d})$.
Appendix B Additional discussions
B.1 Föllmer Flows
B.1.1 Comparison with our work
Instead of implementing a reverse diffusion process, Föllmer Flows Huang et al. [2025], Jiao et al. [2021], Ruzayqat et al. [2023] seek to implement the following Schrödinger bridge:
$\mathrm{d}X_{t}=b(X_{t},t)\mathrm{d}t+\mathrm{d}W_{t},~~X_{0}=0\,$,
with $b(x,t)=\nabla\log\mathbb{E}[f(x+W_{1-t})]$ and where $f$ is given by
$f\coloneqq\frac{\mathrm{d}\mu}{\mathrm{d}\mathcal{N}(0,I_{d})}\,$,
with $\mu$ the target density; for such a process, we have that $X_{1}\sim\mu$. We refer the reader to the notations of Proposition 14 to observe that the shift $b(x,t)$ is given by
$b(x,t)$ $=\nabla\log Q_{-\log(t)/2}(x/\sqrt{t})\,$,
$=\frac{1}{\sqrt{t}}(\nabla\log p_{-\log(t)/2}^{V}(x/\sqrt{t})-\nabla\log\pi(x))\,$,
with $\pi$ is the density of the standard Gaussian. Hence, up to the Gaussian term and a time-rescaling, the shift is exactly given by the intermediate scores of the reverse diffusion. However, because of the Gaussian correction, an extra exponential term appears in the associated self-normalized estimator of the shift: in Huang et al. [2025], Jiao et al. [2021], Ruzayqat et al. [2023], the authors implement
$\hat{b}(x,t)=\frac{\sum_{i=1}^{n}(x+\sqrt{1-t}z_{i}-\nabla V(x+\sqrt{1-t}z_{i}))e^{-V(x+\sqrt{1-t}z_{i})+\|x+\sqrt{1-t}z_{i}\|^{2}/2}}{\sum_{i=1}^{n}e^{-V(x+\sqrt{1-t}z_{i})+\|x+\sqrt{1-t}z_{i}\|^{2}/2}}\,$,
with $z_{i}$ $n$-i.i.d. samples from the standard Gaussian distribution. We suspect that because of the extra exponential terms $e^{\|x+\sqrt{1-t}z_{i}\|^{2}/2}$, as discussed below, their estimator provide appealing guarantees only under very stringent assumptions.
B.1.2 Theoretical guarantees
The theoretical complexity bounds derived in the works of Huang et al. [2025], Jiao et al. [2021], Ruzayqat et al. [2023] quantitatively rely on the assumptions that $f=\frac{\mathrm{d}\mu}{\mathrm{d}\mathcal{N}(0,I_{d})}$ is Lipschitz, has Lipschitz gradient and is bounded from below. In particular, assume for instance that $\mu$ is a standard Gaussian centered at some point $c\in\mathbb{R}^{d}$. In this case, the ratio $f$ reads
$f(x)=e^{2x^{\top}c+\|c\|^{2}}\,$,
which verifies none of the assumptions above if $c\neq 0$. More broadly, even if $f$ verifies these assumptions, the resulting quantities degrade exponentially with the mismatch between $x\mapsto-\log(\mu)(x)$ and $x\mapsto\|x\|^{2}/2$. As noted in Vargas et al. [2022], this limitation is not only a theoretical artifact; in practice, these methods are very unstable and fail to convergence on simple examples.
B.2 Details on the non-smooth example
Consider the mixture $\mu=0.5\mathcal{N}(0,\Sigma_{1})+0.5\mathcal{N}(0,\Sigma_{2})$ with $\Sigma_{1}=diag(1,0.5)$ and $\Sigma_{1}=diag(0.5,1)$. For $(x,y)\in\mathbb{R}^{2}$, it holds that
$\nabla\log(\mu)(x,y)=\frac{(x,2y)e^{-x^{2}/2-y^{2}}+(2x,y)e^{-x^{2}-y^{2}/2}}{e^{-x^{2}/2-y^{2}}+e^{-x^{2}-y^{2}/2}}\,.$
Hence, when $x=y$, we have $\nabla\log(\mu)(x,y)=3/2(x,x)$. Now for $y=x+\eta$, the score reads
$-\nabla\log(\mu)(x,x+\eta)$ $=\frac{(x,2x+2\eta)e^{-x^{2}/2-x^{2}-2\eta x-\eta^{2}}+(2x,x+\eta)e^{-x^{2}-x^{2}/2-\eta x-\eta^{2}/2}}{e^{-x^{2}/2-x^{2}-2\eta x-\eta^{2}}+e^{-x^{2}-x^{2}/2-\eta x-\eta^{2}/2}}$
$=\frac{(x,2x+2\eta)e^{-\eta x-\eta^{2}/2}+(2x,x+\eta)}{e^{-\eta x-\eta^{2}/2}+1}$
$\sim_{x\to+\infty}(2x,x)\,.$
In particular, $\nabla\log(\mu)$ is not Hölder.
B.3 Proof of Proposition 2
Recall we denote $\lambda_{I_{d}n}:=I_{d}n_{i}\lambda_{I_{d}n}(\Sigma_{i}),\quad\lambda_{\max}:=\max_{i}\lambda_{\max}(\Sigma_{i})$. We write:
$\mu(x)=\sum_{i=1}^{p}\tilde{w}_{i}\exp\left(-\frac{1}{2}(x-\mu_{i})^{\top}\Sigma_{i}^{-1}(x-\mu_{i})\right):=\sum_{i=1}^{p}\tilde{w}_{i}\phi_{i}(x)$
with $\tilde{w}_{i}=w_{i}(2\pi)^{-d/2}\det(\Sigma_{i})^{-1/2}$.
Bound on the Hessian.
The Hessian of $\mu$ writes:
$\nabla^{2}\log\mu(x)=\frac{1}{\mu(x)}\nabla^{2}\mu(x)-\frac{1}{\mu(x)^{2}}\nabla\mu(x)\nabla\mu(x)^{\top}$
Since
$\nabla\phi_{i}(x)$ $=-\Sigma_{i}^{-1}(x-\mu_{i})\,\phi_{i}(x)$
$\nabla^{2}\phi_{i}(x)$ $=\left[\Sigma_{i}^{-1}(x-\mu_{i})(x-\mu_{i})^{\top}\Sigma_{i}^{-1}-\Sigma_{i}^{-1}\right]\phi_{i}(x)$
We have
$\nabla\mu(x)$ $=\sum_{i}\tilde{w}_{i}\nabla\phi_{i}(x)=-\sum_{i}\tilde{w}_{i}\phi_{i}(x)\Sigma_{i}^{-1}(x-\mu_{i})$
$\nabla^{2}\mu(x)$ $=\sum_{i=1}^{p}\tilde{w}_{i}\nabla^{2}\phi_{i}(x)=\sum_{i=1}^{p}\tilde{w}_{i}\phi_{i}(x)\left[\Sigma_{i}^{-1}(x-\mu_{i})(x-\mu_{i})^{\top}\Sigma_{i}^{-1}-\Sigma_{i}^{-1}\right]$
Denoting $\gamma_{i}(x):=\frac{\tilde{w}_{i}\phi_{i}(x)}{\mu(x)}\quad$ and $s_{i}(x)=-\Sigma_{i}^{-1}(x-\mu_{i})$ we get
$\nabla^{2}\log\mu(x)$ $=\sum_{i}\gamma_{i}(x)\left[s_{i}(x)s_{i}(x)^{\top}-\Sigma_{i}^{-1}\right]-\left(\sum_{i}\gamma_{i}(x)s_{i}(x)\right)\left(\sum_{j}\gamma_{j}(x)s_{j}(x)\right)^{\top}$
$=\operatorname{Cov}_{\gamma(x)}[s_{i}(x)]-\sum_{i}\gamma_{i}(x)\Sigma_{i}^{-1}$
where the first term is a covariance matrix of the vectors $s_{i}(x)$ under the weights $\gamma_{i}(x)$, hence which is a positive semi-definite matrix. Therefore:
$-\nabla^{2}\log\mu(x)$ $\preceq\sum_{i}\gamma_{i}(x)\Sigma_{i}^{-1}\preceq\sum_{i}\gamma_{i}(x)\cdot\frac{I_{d}.}{\lambda_{I_{d}n}}=\frac{I_{d}.}{\lambda_{I_{d}n}}.$
Bound on the drift.
For the (negative) score of the mixture we have:
$\nabla\log\mu(x)=\sum_{i}\gamma_{i}(x)\Sigma_{i}^{-1}(x-\mu_{i})$
Hence
$\langle-\nabla\log\mu(x),x\rangle$ $=\sum_{i=1}^{p}\gamma_{i}(x)\langle\Sigma_{i}^{-1}(x-\mu_{i}),x\rangle$
$=\sum_{i=1}^{p}\gamma_{i}(x)\left(x^{\top}\Sigma_{i}^{-1}x-\mu_{i}^{\top}\Sigma_{i}^{-1}x\right)\geq\sum_{i=1}^{p}\gamma_{i}(x)\left(\frac{\|x\|^{2}}{\lambda_{\max}}-\mu_{i}^{\top}\Sigma_{i}^{-1}x\right),$
where we have used $\Sigma_{i}^{-1}\succeq\frac{1}{\lambda_{\max}}I_{d}$ for the first term. Now for the second term, using Cauchy-Schwartz and Young’s inequality:
$\mu_{i}^{\top}\Sigma_{i}^{-1}x\leq\|\mu_{i}\|\cdot\|\Sigma_{i}^{-1}x\|\leq\frac{\|\mu_{i}\|\|x\|}{\lambda_{I_{d}n}}\leq\frac{1}{2\lambda_{I_{d}n}}(\|x\|^{2}+\|\mu_{i}\|^{2}).$
Hence, using that $\sum_{i}\gamma_{i}(x)=1$ and that $\sum_{i}\gamma_{i}(x)\|\mu_{i}\|^{2}\leq\max_{i}\|\mu_{i}\|^{2}$ we have:
$\langle-\nabla\log\mu(x),x\rangle$ $\geq\sum_{i=1}^{p}\gamma_{i}(x)\left(\frac{\|x\|^{2}}{\lambda_{\max}}-\frac{1}{2\lambda_{I_{d}n}}(\|x\|^{2}+\|\mu_{i}\|^{2})\right)$
$=\left(\frac{1}{\lambda_{\max}}-\frac{1}{2\lambda_{I_{d}n}}\right)\|x\|^{2}-\frac{1}{2\lambda_{I_{d}n}}\sum_{i=1}^{p}\gamma_{i}(x)\|\mu_{i}\|^{2}.$
$\geq\left(\frac{1}{\lambda_{\max}}-\frac{1}{2\lambda_{I_{d}n}}\right)\|x\|^{2}-\frac{1}{2\lambda_{I_{d}n}}\max_{i}\|\mu_{i}\|^{2}$
$\geq\frac{\|x\|^{2}}{2\lambda_{\max}}-\lambda_{\max}\cdot\max_{i}\left(\frac{\|\mu_{i}\|}{\lambda_{I_{d}n}}\right)^{2}.$
B.4 Discussion on  Lytras and Mertikopoulos [2025]
To justify our claims at the end of Section˜3.2, we first need some preliminary background on Poincaré constants. Let $q\in\mathcal{P}_{\rm ac}(\mathbb{R}^{d})$. We say that $q$ satisfies the Poincaré inequality with constant $C_{P}\geq 0$ if for all $f\in W_{0}^{1,2}(q)$ (functions with zero-mean with respect to $q$ and whose gradient is squared integrable with respect to $q$):
$\int f^{2}(x)dq(x)\leqslant C_{P}\|\nabla f\|^{2}_{L^{2}(q)},$ (21)
and let $C_{P}(q)$ be the best constant in Eq. 21, or $\infty$ if it does not exist. We show that in the case where $\mu=1/2\mathcal{N}(c,\lambda_{max})+1/2\mathcal{N}(-c,\lambda_{min})$ the Poincaré constant is at least $2e^{\frac{\|c\|^{2}}{2\lambda_{max}}}/\|c\|^{2}$ therefore the resulting sampling complexity in Lytras and Mertikopoulos [2025, Theorem 3] is at least $O(\text{poly}(\|c\|^{2}e^{\frac{\|c\|^{2}}{2\lambda_{max}}},\epsilon^{-1}))$.